# Distinguishing Ignorance from Error in LLM Hallucinations

### v1

@@ Abstract
Large language models (LLMs) are susceptible to hallucinations—outputs that are ungrounded, factually incorrect, or inconsistent with prior generations. We focus on close-book Question Answering (CBQA), where previous work has not fully addressed the distinction between two possible kinds of hallucinations, namely, whether the model (1) does not hold the correct answer in its parameters or (2) answers incorrectly despite having the required knowledge. We argue that distinguishing these cases is crucial for detecting and mitigating hallucinations. Specifically, case (2) may be mitigated by intervening in the model’s internal computation, as the knowledge resides within the model’s parameters. In contrast, in case (1) there is no parametric knowledge to leverage for mitigation, so it should be addressed by resorting to an external knowledge source or abstaining. To help distinguish between the two cases, we introduce Wrong Answer despite having Correct Knowledge (WACK), an approach for constructing model-specific datasets for the second hallucination type. Our probing experiments indicate that the two kinds of hallucinations are represented differently in the model’s inner states. Next, we show that datasets constructed using WACK exhibit variations across models, demonstrating that even when models share knowledge of certain facts, they still vary in the specific examples that lead to hallucinations. Finally, we show that training a probe on our WACK datasets leads to better hallucination detection of case (2) hallucinations than using the common generic one-size-fits-all datasets.  [^1] Code and datasets at https://github.com/technion-cs-nlp/hallucination-mitigation.
@@ 1 Introduction
Large Language Models (LLMs) are prone to generating outputs that are not grounded in the model’s input or real-world facts, as well as outputs that may be inconsistent with earlier generations within the same session [Ji et al., 2023, Sharma et al., 2023, Kalai and Vempala, 2023]. These issues are collectively referred to as hallucinations, and they are crucial to address since they reduce the reliability of LLMs.
Numerous studies have focused on the detection and mitigation of hallucinations (e.g. Li et al. [2023], Zhang et al. [2024], Marks and Tegmark [2023], Chen et al. [2024], CH-Wang et al. [2023]). However, existing work often fails to distinguish between the different causes of hallucinations, conflating two distinct types: first type, denoted as HK-, refers to cases where the model lacks the required information, leading it to hallucinate. The second, denoted as HK+, type occurs when, although the model has the necessary knowledge and can generate correct answers under certain prompts, it still produces an incorrect response in a different but similar prompt setting. These types represent fundamentally different problems, requiring different solutions: When a model lacks knowledge one should consult external sources (or abstain), but when a model has the knowledge it may be possible to intervene in its computation to obtain the correct answer. Failing to differentiate between these causes can weaken the effectiveness of detection and mitigation techniques, which often categorize outputs simply as either ’hallucination’ or ’factually correct’ without further investigating these two distinct types of hallucination [Marks and Tegmark, 2023, Azaria and Mitchell, 2023, Li et al., 2023, Rateike et al., 2023, Zhang et al., 2024, Chen et al., 2024, Zou et al., 2023, Hoscilowicz et al., 2024].
Our first contribution in this work is an automatic approach to obtain Wrong Answers despite having Correct Knowledge (WACK). This approach constructs a model-specific hallucination dataset that captures the distinction between the two types of hallucinations. It differentiates between hallucinations caused by a lack of knowledge (HK-) and those caused by incorrect generation despite the existence of the knowledge in the target model (HK+). The process infers by automatically categorizing examples based on the model’s knowledge via inspection of correct responses in output samples. If the model lacks the required knowledge, the example is labeled as HK-. Otherwise, for examples where the model possesses the correct knowledge, it further splits the cases into “factually-correct” and HK+ based on the model generation in an altered prompt setting. This alternate prompt setting employs persuasion [Xu et al., 2023, Zeng et al., 2024], weak semantics [Yao et al., 2023] that aim to modify the prompt to reduce its semantic content, and other techniques to induce hallucinations in scenarios that mimic regular interactions with a model. The WACK method is illustrated in Figure 1.
We construct WACK datasets for three state-of-the-art LLMs of the 7B–9B size range. With these tailored benchmarks, we investigate how different types of hallucinations (HK- and HK+) are represented within the models. We do so by training probes on the model’s inner states, a common method in the field of hallucination detection [CH-Wang et al., 2023, Azaria and Mitchell, 2023]. While prior work detects whether any hallucination occurs, we show it is possible to distinguish between the two hallucination types (Section 4.1). Next, we focus on HK+ type of hallucinations and show the generalization of WACK between different prompt settings; a probe trained on examples from one setting is able to predict hallucinations in another setting (Section 4.2). Lastly, we show that WACK datasets differ across models, highlighting the significance of using model-specific datasets that account for each model’s unique knowledge and hallucination patterns (Section 5.1). As a result, we demonstrate that model-specific datasets are more effective for HK+ detection than generic datasets (Sections 5.2 and 5.3).
Our main contributions are as follows: (I) We propose WACK, a methodology for constructing a model-specific dataset that includes factually correct and hallucination examples both due to lack of knowledge (HK-) and despite having knowledge (HK+). We will release the datasets for the models we experimented with. (II) We demonstrate that a model’s internal states can be used to distinguish between the two hallucination types. (III) We demonstrate the importance of model-specific datasets for HK+ hallucination detection.
@@
Figure 1: WACK Setup: (a) The first step in our process involves detecting whether the model knows the correct answer. If the model does not know the correct answer, the example is labeled as hallucination caused by not knowing (HK-) If the model knows the correct answer, we proceed to the next stage. (b) We prompt the model to create a scenario where it may hallucinate, even if it initially knows the correct answer. Here we show a snowballing bad-shots prompt. (c) Under the new setting, if the model generates the correct answer, the example is labeled as factually-correct; otherwise, it is labeled as a hallucination despite knowing (HK+).
2 Model-specific dataset Construction
@@
In this section, we outline the process of creating a model-specific dataset, with a focus on generating HK+ hallucination examples. Figure 1 provides a detailed overview of this setup. The process begins by classifying examples based on the model’s knowledge, labeling all instances where the model lacks knowledge as HK-. Next, using examples where the model knows the correct answer, we create a scenario in which hallucinations can occur despite the model’s knowledge (HK+). The next two subsections describe these steps. We focus on closed-book question answering (CBQA) tasks with short answers.
@@ 2.1 Categorization of Knowledge
In our CBQA setting, a model is given a question $q$ and generates an answer $\tilde{a}$, which may match the factually correct gold answer $a_{g}$ or else constitute a hallucination. Knowledge in a language model can be viewed as lying on a spectrum. At the low-knowledge end, the knowledge stored in the model’s parametric has little to no association between $a_{g}$ and $q$, while at the high-knowledge end, there is a high association. Hallucinations at the low-knowledge end of the spectrum are somewhat expected, as the model is unlikely to generate $a_{g}$ (that is, we expect $\tilde{a}\neq a_{g}$). However, hallucinations can occur anywhere along this spectrum, including at the high-knowledge end. Detecting the cause of hallucinations in the middle of the spectrum is more complex, as they may arise either from insufficient knowledge or despite adequate knowledge.
@@
To simplify our analysis, we focus on the two ends of the spectrum: high-knowledge and low-knowledge, which still provide a compelling overview of the two types of hallucinations. To this end, given model $M$, we follow the setup of Gekhman et al. [2024] in which $M$ generates various completions to $q$, and then we verify the existence of the answer $a_{g}$ in the output. Specifically [^2] The decision process of the following hyperparams can be found in Appendix A., we perform one greedy generation plus five generations with a temperature of $0.5$. We use a 3-shots in-context learning scenario [Brown et al., 2020], generate a maximum of 5 tokens, and look for an exact match to $a_{g}$. If the model did not generate $a_{g}$ in any of the generations, the example is labeled HK-. If the model generates $a_{g}$ in all the attempts, this example is considered a high-knowledge scenario and we next label it as either factually-correct or HK+.
2.2 Hallucination Despite knowledge
To label a high-knowledge example as either factually correct or HK+, we follow Zhang et al. [2023], who demonstrated that after a model produced an incorrect answer, it was likely to generate an incorrect explanation to justify its error, which they termed the “snowballing effect”. Similar behaviors were also shown when using persuasion techniques that modify the prompt to include persuasions [Xu et al., 2023, Zeng et al., 2024], and weak semantics [Yao et al., 2023] that modify the prompt to reduce its semantic content.
We argue that these settings are important to focus on as they reflect mistakes in the prompt, which can originate from either the user or the model’s previous generation.
  1. 1.
    User mistakes: We cannot expect users to have perfect knowledge—they can write a wrong fact or make grammar or language mistakes—and if their own errors cause the model to hallucinate, this represents a real-world problem that needs to be addressed.
  2. 2.
    Model’s previous mistakes: The model may create a snowballing effect on its own by generating mistakes in previous turns that would be added to the prompt (e.g., Zhang et al. 2023). If the model produces a hallucination, we do not want this to affect subsequent generations, thus this issue should also be mitigated.
To facilitate the creation of HK+ hallucinations in large quantities, we design two synthetic setups that align with the ideas described above: (a) Bad-shots, leveraging snowballing, and (b) Alice-Bob, leveraging persuasion and weak semantics. Later in the paper, we demonstrate that one setting generalizes to the other one, indicating that it is valid to use these specific setups to investigate the general phenomenon of HK+.
Bad-shots setting.
This setting illustrates how mistakes in the context can create a cascading effect that may compromise the accuracy of the model’s subsequent generations. To imitate the snowballing effect on a large scale, we propose a synthetic method which we name Bad-Shot Prompting. We construct 20 false QA pairs using ChatGPT 3.5 [OpenAI, 2022], where the false answer is semantically similar to the correct one. For instance, here is a good-shot example and its corresponding bad-shot example:
missingfnum@@desciitemGood-shot
question: Which element has the chemical symbol ’H’? answer: Hydrogen
missingfnum@@desciitemBad-shot
question: Which element has the chemical symbol ’H’? answer: Helium
For each target question in the eval set we prepend bad-shot examples in a few-shot manner, thus simulating mistakes that a user or a model might create as part of the input context. In practice, we use 3 random bad-shots before each example from the dataset, as demonstrated in Figure 2.
question: In what year did World War II end? answer: 1939 question: What is the smallest prime number? answer: 1 question: Who wrote ’Romeo and Juliet’? answer: Jane Austen question:What does the USA celebrate on the first Monday in September? answer:
Figure 2: Example prompt using 3-bad-shot snowballing. Depending on the model’s response, this example would be labeled as either ‘factually correct’ or ‘hallucinated’ (HK+).
@@ 2.2 Hallucination Despite knowledge
As we will see below, the bad-shot setting enables us to obtain many examples in which models hallucinate despite having the knowledge to answer correcly (HK+). In Table 1 we show one example per model from the TriviaQA dataset [Joshi et al., 2017], where the model generated the correct answer using 3 random good shots (regular-few-shots), but hallucinated when prompted with 3 random bad shots.
Table 1: Generated answers using greedy decoding on TriviaQA.
Generation
Model Prompt w/ good-shots w/ bad-shots
Gemma-2-9B question: What does the USA celebrate on the first Monday in September? answer: Labor day Independence Day
Llama-3.1-8B question: In which Disney film do the fairies Flora, Fauna and Merryweather appear? answer: Sleeping Beauty The Lion King
Mistral-7B-v0.3 question: What is the official spoken language of China? answer: Mandarin English
@@
The Alice-Bob setting uses persuasion and weak semantics in the form instead of snowballing bad shots. In this setting, we add the following text at the beginning of the prompt, along with a one-shot example at the end:
@@
This setting is more subtle than the bad-shots setting and generates fewer hallucinations. Drawing on ideas from related studies [Xu et al., 2023, Yao et al., 2023, Zeng et al., 2024], the aim is to simulate a persuasion scenario with a few deliberate mistakes (underlined in the prompt). The prompt typos are intended to mimic small inaccuracies that simulate an error. The persuasive aspect of the setting comes through several nuances in the text: (1) there is an implication that Bob is not smart, (2) the test is described as difficult, (3) to pass, one only needs to be correct on 2 out of the 4 questions, and (4) there is no suggestion that exceeding the minimum required score offers any advantage.
2.3 Dataset Construction
Equipped with our process for separating examples of low and high knowledge (Section 2.1) and and further labeling high-knowledge examples (Section 2.2), we create model-specific datasets. As sources for examples to label, we use two common closed-book question answering datasets: TriviaQA [Joshi et al., 2017] and NaturalQuetions [Kwiatkowski et al., 2019]. We experiment with three models: Mistral-7B-v0.3 [Jiang et al., 2023], Llama-3.1-8B [Dubey et al., 2024] and Gemma-2-9B [Team et al., 2024].
Table 2 provides the number of examples in each category for the resulting model-specific datasets. We observe that even under the bad-shots setting, most of the model’s high-knowledge examples are labeled as factually correct rather than hallucinations. Still, we are left with sufficient cases of hallucinations-despite-knowledge (HK+), which we use in the subsequent sections. In the Alice-Bob setting, we observe similar trends, but with fewer hallucinations (Appendix C). For more details regarding the dataset construction, see Appendix B.
Table 2: Dataset labels statistics using bad-shot setting.
# Factually # Hallucination # Do-not-know
Dataset correct (HK+) (HK-)
TriviaQA-Llama3-WACK 14154 1675 7356
Natural-Questions-Llama3-WACK 5934 1104 14739
TriviaQA-Gemma-WACK 13534 2563 6991
Natural-Questions-Gemma-WACK 6045 1859 13762
TriviaQA-Mistral-WACK 12652 2841 7650
Natural-Questions-Msitral-WACK 5562 1546 14689
3 Implementation details
We aim to show the importance of separating the two hallucination types and using our model’s specific dataset to create better detectors. In the following sections, we report on various experiments for detecting different types of hallucinations by training classifiers on inner model states. In all detection experiments, we randomly select 1000 examples from each label for analysis in each dataset and split them to 70%/30% for training/test. [^3] In cases where there are fewer than 1000 hallucinations we use all the hallucinations we have. We use a linear classifier for detection, as in prior work [Li et al., 2023, CH-Wang et al., 2023]. [^4] We ran the detection on normalized vectors of the model’s inner states at the last token using a linear SVM. The detection results in the main paper are on hidden states from the residual component (after each Transformer block); see Appendix D for similar results on the MLP and Attention components. Each experiment was repeated with three random seeds for the SVM and split into training/test sets. We report average results with standard deviations. To maintain consistency with the prompts used in the creation of the ‘hallucination despite knowledge’ dataset setting, all examples incorporate similar prompts (bad shots or Alice-Bob). In addition, unless stated otherwise the results are shown under the bad-shot setting. Lastly, unless stated otherwise, we use examples that include the concatenated model-generated answer (which may or may not be a hallucination) along with the prompt setup and question.
All experiments were run on NVIDIA RTX 6000 Ada (49GB) with 4 CPUs. Generating all the datasets and results takes approximately 2 weeks on one GPU.
4 Detecting Different Types of Hallucinations
This section first investigates whether different types of hallucinations are represented differently inside models. Then we examine the relationship between different settings used for inducing hallucinations.
4.1 Models can Distinguish Hallucinations despite knowledge from hallucinations caused by not knowing from factually correct
We first explore the distinction between hallucinations arising from a model’s lack of knowledge and those that occur even when the model possesses relevant information. This differentiation is crucial for understanding hallucinations’ underlying mechanisms and developing targeted detection and mitigation strategies. We employ detection from the model’s inner state to demonstrate that the model represents these hallucinations differently. Note that this is a challenging task, as it not only requires distinguishing hallucinations from factually correct responses, but also determining the type of hallucination. This involves an understanding of the model’s knowledge, which must be considered.
To address a comprehensive scenario, we differentiate between cases where the model: (1) knows the information and does not hallucinate (’factually correct’), (2) knows the information but hallucinates (’HK+’), and (3) does not know the information and thus hallucinates (’HK-’). In Figure 3 (Top), we see the detection accuracy results for the 3 classes across the model’s layers using the bad-shot setting. The accuracy at the highest layer is 60%–70%, well above the random baseline of 33%. Additionally, when we evaluate the accuracy of any two of the three classes produced by the classifier in the middle layer (16), we observe that the accuracy is no less than 70%. This high result indicates that models’ inner states contain information for differentiating between the 3 cases. Figure 3 (Bottom) shows similar trends using the Alice-Bob setting, although the results are lower in this case, as may be expected given the subtlety of this setting.
(a) TriviaQA, bad-shots.
(b) Natural Questions, bad-shots.
(c) TriviaQA, Alice-Bob.
(d) Natural Questions, Alice-Bob.

### v2

@@ Abstract
Large language models (LLMs) are susceptible to hallucinations—factually incorrect outputs—leading to a large body of work on detecting and mitigating such cases. We argue that it is important to distinguish between two types of hallucinations: ones where the model does not hold the correct answer in its parameters, which we term HK-, and ones where the model answers incorrectly despite having the required knowledge, termed HK+. We first find that HK+ hallucinations are prevalent and occur across models and datasets. Then, we demonstrate that distinguishing between these two cases is beneficial for mitigating hallucinations. Importantly, we show that different models hallucinate on different examples, which motivates constructing model-specific hallucination datasets for training detectors. Overall, our findings draw attention to classifying types of hallucinations and provide means to handle them more effectively.  [^1] Code and datasets at https://github.com/technion-cs-nlp/hallucination-mitigation.
Distinguishing Ignorance from Error in LLM Hallucinations
Adi Simhi1     Jonathan Herzig2     Idan Szpektor2     Yonatan Belinkov1 thanks: adi.simhi@campus.technion.ac.il 1Technion – Israel Institute of Technology   2Google Research
@@ 1 Introduction
LLMs are prone to generate outputs that lack grounding in the model’s input or in real-world facts, as well as outputs that may be inconsistent with earlier generations within the same session (Ji et al., 2023; Sharma et al., 2023; Kalai and Vempala, 2023). These issues, collectively known as hallucinations, are critical to address due to their impact on LLM reliability.
Numerous studies have focused on the detection and mitigation of hallucinations (e.g. Li et al., 2023; Zhang et al., 2024a; Marks and Tegmark, 2023; Chen et al., 2024; CH-Wang et al., 2023). However, existing work does not fully investigate the different causes of hallucinations, often blending two distinct types:
HK-
The model lacks the required information for a correct answer, leading it to hallucinate.
HK+
Although the model generates correct answers under certain prompts, it still produces incorrect responses in different but similar prompt settings.
These types represent fundamentally different problems, requiring different solutions: When a model lacks knowledge, it should consult external sources or abstain, but when a model inherently holds the knowledge in its parameters, it may be possible to prompt it or intervene in its computation to obtain the correct answer.
@@
Figure 1: WACK Setup: (a) The first step in our process involves detecting whether the model knows the correct answer. If the model does not know the correct answer, the example is labeled as hallucination caused by not knowing (HK-) If the model knows the correct answer, we proceed to the next stage. (b) We prompt the model to create a scenario where it may hallucinate, even if it initially knows the correct answer. Here we show a snowballing prompt. (c) Under the new setting, if the model generates the correct answer, the example is labeled as factually-correct; otherwise, it is labeled as a hallucination despite knowing (HK+).
In this work, we demonstrate the prevalence of HK+, showing that it occurs in up to 24% of cases (Section 3.3). Importantly, we observe that HK+ errors manifest even when using mild prompts that neither contain misinformation nor induce hallucination-prone scenarios. Thus, these hallucinations are not merely a consequence of flawed human prompts but can arise independently.
To study HK+, HK-, and factually-correct examples, we developed an automatic approach to collect Wrong Answers despite having Correct Knowledge (WACK). Our approach (Figure 1) constructs a model-specific hallucination dataset that labeled hallucination types —HK- and HK+— by introducing novel ways of constructing HK+. The process assesses the target model’s knowledge by inspecting the number of correct responses in a neutral prompt out of multiple output samples for questions with curated known answers. For questions in which the model indicates knowledge, WACK explores altered prompt settings, which induce hallucinations in scenarios that mimic typical interactions with an LLM. These prompts reveal HK+ examples, in which the model answers incorrectly even though it holds the knowledge for a correct answer.
After establishing the prevalence of HK+ and creating a framework to investigate the hallucination cases, we provide a proof of concept showing that mitigation can benefit from distinguishing HK+ from HK- (Section 4.1). Next, we investigate how different types of hallucinations (HK- and HK+) are represented within models. We do this by training linear classifiers on each model’s internal states, a common technique in hallucination detection research (CH-Wang et al., 2023; Azaria and Mitchell, 2023). While prior work focuses on detecting whether any hallucination occurs, we demonstrate that simple classifiers can distinguish between the two hallucination types to some extent (Section 4.2). To further demonstrate the utility of our setup for investigating HK+, we show that HK+ generalizes across different prompt settings; a classifier trained on examples from one prompting setting is able to predict hallucinations in another setting (Section 5.2).
Our last contribution shows that the WACK datasets differ between models, highlighting the importance of using model-specific datasets that account for the unique knowledge and hallucination patterns of each model (Section 5.1). As a result, we demonstrate that model-specific datasets are more effective for HK+ detection than datasets that ignore the HK-/HK+ distinction (Sections 6.1 and 6.2). Using this novel framework and its results, we can improve the detection and mitigation of hallucinations.
Our main contributions are: (I) Indicate the prevalence of HK+ hallucinations by showing their consistency under varying prompts, including supposedly factually-encouraging prompts. (II) Demonstrating that mitigation can benefit from the separation of the two types of hallucination. (III) Proposing WACK, an approach for constructing model-specific datasets that include factually correct and hallucination examples both due to lack of knowledge (HK-) and despite having knowledge (HK+). We will release the datasets we constructed. (IV) Showing the importance of model-specific datasets for HK+ detection and the ability for preemptive detection.
2 Related Work
Response
Model Prompt Original Truthful setting
Gemma-2-9B question: What is the technical term for a solid figure with five plane (flat) faces? answer: pentahedron pentagon
Llama-3.1-8B question: What is the national animal of Greenland? answer: Polar bear Arctic fox
Mistral-7B-v0.3 question: The Java Trench is part of which ocean?answer: Indian Ocean Pacific Ocean
Table 1: Generated answers using greedy decoding on TriviaQA, illustrating the HK+ using Truthful setting.
Hallucinations.
One approach for detecting hallucinations treats the model as a black box, posing questions or sampling its outputs (Gekhman et al., 2023; Pacchiardi et al., 2023; Manakul et al., 2023; Li et al., 2024a). Another line of works examines the model’s hidden representations, often via a classifier (Burns et al., 2022; He et al., 2023; Rateike et al., 2023; Slobodkin et al., 2023; Azaria and Mitchell, 2023; CH-Wang et al., 2023; Yuksekgonul et al., 2023; Chen et al., 2023; Yin et al., 2024; Levinstein and Herrmann, 2024; Marks and Tegmark, 2023; Li et al., 2023). There is also ongoing research on mitigating hallucinations by steering the model towards correct answers, including prompting (Si et al., 2022; Dhuliawala et al., 2023), fine-tuning (Tian et al., 2023; Razumovskaia et al., 2023), and modifying model logits (Chuang et al., 2023; Kalai and Vempala, 2023; Li et al., 2023; Zhang et al., 2024a). These works do not distinguish between types of hallucinations, which may hinder performance of both detection and mitigation. Specifically, in Section 4.1 we show that mitigation via prompting is plausible for HK+ but not for HK-, and that separating between HK+ and HK- would result in performance improvement.
Most prior work in this area uses generic datasets. Some works explore model-specific hallucination datasets and demonstrated their importance (Azaria and Mitchell, 2023; Ji et al., 2024; Cao et al., 2023; CH-Wang et al., 2023). However, they do not distinguish between the causes of hallucinations. We show that model-specific datasets with HK- and HK+ typing can be beneficial for detection and mitigation, and that they are useful for preemptive detection (Section 6).
Knowledge.
Several studies attempt to assess the knowledge stored in a model after it has been trained, including prompting, sampling, using paraphrases, or training to select model knowledge (Petroni et al., 2019; Jiang et al., 2020; Elazar et al., 2021; Kadavath et al., 2022; Gekhman et al., 2024; Jiang et al., 2024; Ren et al., 2023; Zhang et al., 2024b). Similarly in this work, we use greedy decoding and sampling to evaluate the model’s knowledge.
A few recent studies detect cases of missing knowledge (Wen et al., 2024; Zhang et al., 2024b), similar to our HK- hallucinations. However, they focus on using external knowledge to mitigate them while we focus on the advantage of HK- and HK+ typing in understanding the model’s activations.
Jailbreaking.
Jailbreaking refers to techniques that cause LLMs to generate unexpected or incorrect answers. For instance, Zhang et al. (2023) demonstrate the snowballing effect, where once the model outputs a wrong fact, it is more likely to generate an incorrect explanation for that fact. Additionally, research found that a model’s answers can change due to persuasion, long conversations, fantasy settings, LLM personas, and out-of-distribution prompts (Zeng et al., 2024; Li et al., 2024b; Xu et al., 2023; Yao et al., 2023; Nardo, 2023; Joshi et al., 2023; Pacchiardi et al., 2023; Meinke et al., 2024). These studies show that prompt characteristics heavily influence output correctness, causing hallucinations even when the model knows the answer (HK+).
Yet, they only demonstrate the phenomenon but do not explore its representation, detection, and mitigation. In this work we dive into the HK+ phenomenon, define it properly, quantify its occurrence across models, prompts and datasets, and emphasize its importance when attempting to mitigate hallucinations.
3 WACK Framework
@@
In this section, we outline the creation of model-specific datasets for the closed-book question answering (CBQA) setting with short answers, with a focus on generating HK+ examples. Figure 1 provides an overview of our approach. The first step classifies examples based on the model’s knowledge, labeling all instances where the model lacks knowledge as HK-. Next, for each example where the model has the knowledge for a correct answer, we create a prompting scenario in which hallucinations could occur despite the model’s knowledge (HK+). The next subsections describe these steps.
TriviaQA Natural Questions
Truthful Persona Alice-Bob Snowballing Truthful Persona Alice-Bob Snowballing
Gemma 3.97% 4.05% 4.22% 15.92% 8.24% 8.97% 8.98% 23.52%
Llama 5.32% 4.85% 6.18% 10.58% 11.86% 12.57% 13.91% 15.69%
Mistral 4.94% 5.93% 6.38% 18.34% 9.67% 12.52% 12.32% 21.75%
Table 2: Percentage of HK+ in TriviaQA and Natural Questions datasets with different settings. As the setting becomes more extreme (moving from left to right in each block), the HK+ percentage increases.
@@ 3.1 Categorization of Knowledge
In our CBQA setting, given a question $q$ with a known gold answer $a_{g}$, a target model generates an answer $\tilde{a}$, which may match $a_{g}$ or else constitute a hallucination. Knowledge in a language model can be viewed as lying on a spectrum. We refer to the model’s parametric knowledge as being at the ‘low-knowledge end’ when there is little to no association between $a_{g}$ and $q$, and at the ‘high-knowledge end’ when this association is strong. The model is unlikely to generate the correct answer in the low-knowledge end, whereas in the high-knowledge end, it typically does.
Hallucinations at the low-knowledge end of the spectrum are somewhat expected, as the model is unlikely to generate $a_{g}$ (that is, we expect $\tilde{a}\neq a_{g}$). However, hallucinations can occur anywhere along this spectrum, including at the high-knowledge end. Detecting the cause of hallucinations in the middle of the spectrum is more complex, and is beyond the scope of the current work.
@@
To simplify our analysis, we focus on the two ends of the spectrum: high-knowledge and low-knowledge, which still provide a compelling overview of the two types of hallucinations. Given model $M$, we follow the setup of Gekhman et al. (2024) in which $M$ generates various completions to $q$, and then we verify the existence of the answer $a_{g}$ in the output. [^2] Specifically, we perform one greedy generation plus five generations with a temperature of $0.5$. We use a 3-shots in-context learning scenario (Brown et al., 2020), generate a maximum of 5 tokens, and look for an exact match to $a_{g}$. See Appendix A for the decision process of the hyperparams. If the model did not generate $a_{g}$ in any of the generations, the example is labeled HK-. If the model generates $a_{g}$ in all attempts, the example is considered a high-knowledge scenario, to be labeled as either factually-correct or HK+.
Our categorization of examples according to the model’s knowledge may not be perfect. However, as Appendix A shows, different configurations (temperature, number of generations, etc.) produce similar labeling outcomes, which gives confidence in the process.
3.2 HK+ Settings
To label a high-knowledge example as either factually correct or hallucination (HK+), we design four synthetic setups that nudge the model toward producing an incorrect answer with progressively stronger settings. These setups are inspired by recent observations that various prompt modifications induce hallucinations, including perturbations (Yao et al., 2023), adding persuasive language (Xu et al., 2023; Zeng et al., 2024), and other “snowballing” effects, where one error by a model makes it generate incorrect explanations to justify the error (Zhang et al., 2023) and similarly that contextual errors can make the model generate more mistakes (Meinke et al., 2024). Notably, we observe that even completely error-free settings can produce HK+ examples. Later in the paper, we demonstrate that one setting generalizes to another, indicating that it is valid to use these synthetic setups to investigate the general phenomenon of HK+.
Truthful setting.
This setting represents a simple prompt modification, which should encourage the model to generate truthful responses. We prompted ChatGPT (OpenAI, 2022) to generate 10 paraphrased variations of the instruction: “Generate answers that are entirely factual and precise”. For each example, we included a one-shot example with a permutation of the text. The 10 paraphrased prompts are provided in Appendix B.2.
@@ 3.2 HK+ Settings
Table 1 shows one example per model from the TriviaQA dataset (Joshi et al., 2017), where the model generated the correct answer using 3 random good shots (regular-few-shots), but hallucinated when prompted with this setting.
Persona setting.
To create a diverse, persona-driven approach to truthful response generation, we selected 10 personas from Ge et al. (2024) and prompted ChatGPT to rewrite the following text as if each persona wrote it: “I would really like to know the answers to the following very hard questions. Could you please explain them to me?”. For each example, we again incorporated a one-shot example, pairing a permutation of the text with a different persona. The 10 rewritten prompts are provided in Appendix B.2.
@@
The Alice-Bob setting uses persuasion and text perturbations (typos) to simulate an error, inspired by related studies (Xu et al., 2023; Yao et al., 2023; Zeng et al., 2024). Specifically, we add the following text at the beginning of the prompt (typos are underlined), along with a one-shot example at the end:
@@
The persuasive aspect of the setting comes through several nuances in the text: (1) there is an implication that Bob is not smart, (2) the test is described as difficult, (3) to pass, one only needs to be correct on 2 out of the 4 questions, and (4) there is no suggestion that exceeding the minimum required score offers any advantage.
Snowballing setting.
This setting illustrates how mistakes in the context can create a cascading effect that may compromise the accuracy of the model’s subsequent generations. Specifically, we construct 20 false QA pairs using ChatGPT 3.5, where the false answer is semantically similar to the correct one (Appendix B.2). For instance, given the question “Which element has the chemical symbol ’H’”, with the gold answer “Hydrogen”, we obtained the answer “Helium”. For each test question, we prepend $k$ random snowballing examples in a few-shot manner, thus simulating mistakes that a user or a model might create as part of the input context. [^3] We use $k=3$ in the main paper. The results are consistent with other choices; Appendix K.
This setting is crucial to study as it reflects prompt diversity or errors, which can stem from either the user or the model’s prior outputs. User errors, such as incorrect facts or language mistakes, represent real-world challenges that can lead to model hallucinations and must be addressed. Similarly, errors in the model’s earlier outputs can create a snowballing effect, compounding mistakes in subsequent prompts and generations. Mitigating both user- and model-originated errors is essential for reducing hallucinations and ensuring reliable performance.
Model Dataset HK+ mitigation (%) HK- mitigation (%)
Gemma TriviaQA 17.2 / 21.4 / 23.4 / 9.0 1.8 / 2.2 / 1.8 / 2.6
Natural Questions 20.4 / 27.4 / 28.0 / 14.6 2.4 / 2.4 / 2.4 / 2.0
Llama TriviaQA 19.4 / 27.0 / 16.6 / 16.6 1.6 / 1.6 / 1.0 / 2.0
Natural Questions 16.6 / 24.4 / 25.2 / 21.0 2.0 / 1.6 / 1.2 / 1.2
Mistral TriviaQA 27.4 / 27.8 / 20.0 / 49.2 1.6 / 3.2 / 0.8 / 2.8
Natural Questions 34.2 / 35.2 / 23.0 / 51.0 2.4 / 3.0 / 1.4 / 2.8
Table 3: Comparison of HK+ and HK- mitigation success percentage results on Truthful / Persona / Alice-Bob / Snowballing settings. Mitigating HK+ is significantly more successful than mitigating HK-.
(a) Gemma.
(b) Llama.
(c) Mistral.